OpenStudy (anonymous):

Algebra 1 help please?? (5.3 x 10^-7)(6 x 10^-5) (x is not a variable it just means to multiply)

OpenStudy (anonymous):

@rosedewittbukater @Squirrels @Whitemonsterbunny17

OpenStudy (anonymous):

@MTALHAHASSAN2 @Mesopretty @MyaDaBoo

OpenStudy (anonymous):

3.18e^-11 Is there answer choices?

OpenStudy (anonymous):

nope :( whats the e?

OpenStudy (anonymous):

forget the e just put 3.18^-11

OpenStudy (anonymous):

don't do this sigh either ^ thats just example

OpenStudy (anonymous):

ok so i forgot i was supposed to put it in standard form :(

OpenStudy (anonymous):

so I got it wrong, but it lets me try again

OpenStudy (anonymous):

So your suppose to put it in standard form?

OpenStudy (anonymous):

yes, but it changed out the numbers

OpenStudy (anonymous):

The new numbers are...

OpenStudy (anonymous):

(3.7 x 10^-4)(3 x 10^-3)

OpenStudy (anonymous):

Can you copy and paste the whole question?

OpenStudy (anonymous):

no, it won't let me:( but i can type it in word for word, hold on a sec

OpenStudy (triciaal):

when you multiply numbers(0r variables) with the same base you add the exponents

OpenStudy (anonymous):

Okay that'll work

OpenStudy (triciaal):

group like terms and simplify

OpenStudy (anonymous):

Evaluate the following expression using exponential rules. Write the result in standard notion.

OpenStudy (triciaal):

to put in scientific notation a* 10^n where a is between 0 and 10 , and n is an integer

OpenStudy (anonymous):

ok so...............................

OpenStudy (anonymous):

um idk

OpenStudy (jdoe0001):

\(\large { \begin{array}{cllll} (5.3\times 10^{-7})(6\times 10^{-5})\\ (5.3\times 10^{-7})(6.0\times 10^{-5}) \end{array} }\) a negative exponent over the 10, means, "move the decimal point to the LEFT that many slots"

OpenStudy (anonymous):

so do i first times 5.3 by 6? Then move the decimal 12 places to the left?

OpenStudy (triciaal):